The Five-Vector Correlation and Neutron Depolarization

PEN collaboration at KEK

November 22, 1994

There had been many experimental attempts to find out an evidence of Time Reversal Non-Conservation (TRNC), but only the upper limit had been obtained. We are going to perform experiments to observe TRNC by measuring spin correlations in neutron-nucleus interactions in which the resonance enhancement is expected as same as the parity nonconservation effect. In the case of parity nonconservation at ^{139}La 0.734eV resonance, the enhancement factor is about 10^{6}. There are two experimental possibilities to measure TRNC using neutron induced resonance reactions. The first one is the measurement of three-vector correlation parameters and the second is the measurement of five-vector correlation parameters. In the former case, Neither parity nor time reversal invariance is conserved , where as in the latter case only the parity is conserved. We focus on the measurement of five-vector correlation parameters in this report. The measurement of the five-vector correlation parameter is performed with an experimental set up shown schematically in Fig 1. Plarized neutrons transmit the aligned nuclei target. The correlation amplitude is proportional to s⃗ · (k⃗ × I⃗)(k⃗ · I⃗), where s⃗ is the neutron spin, k⃗ is the neutron momentum and I⃗ is the spin of target nuclei. It is clear that this term is parity even and time reversal odd. This term can be derived from the measurement of the asymmetry with the neutron spin up and down perpendicular to the plane formed by the neutron momentum and the target alignment axis.

Figure 1: 5 Vector Correlation experiment

218z The nucleus of ^{165}Ho can be aligned relatively easily. ^{165}Ho is ferromagnetic below 20 K and,magnetized to the direction of the C-axis. If external magnetic field is zero a Ho crystal is divided into small domains and the overall crystal magnetization is zero, because there are the same number of domains which are magnetized to opposite directions. At this temperature (20K) only the electrons are aligned. By cooling the crystal down below a few hundred mK the Ho nuclei are also aligned due to the large hyperfine interaction.

A measurement of five-vector correlation terms using a Ho single crystal was performed by Koster et al.[1] with 2 MeV neutrons. They found that the analyzing power of the fivevector correlation is (1 ± 6) × 10^{−4}. However the five-vector correlation experiment at the resonance region of Ho, where the enhancement is expected, has not yet been performed.

In the case of such experiments at the resonance region, the depolarization problem arises. When the polarized neutron transmits through a domain of the Ho crystal, the neutron spin is rotated due to the magnetic field. Since the total magnetization is zero, the averaged total rotation angle should be zero. However the variance of the rotation is not zero, then the neutron polarization decreases after transmitting through a crystal. Therefore it is necessary to measure the neutron depolarization in a Ho crystal in the resonance region. Haase et al.[2](a LosAlamos Group) measured the depolarization at 0.734eV and found that the polarization perpendicular to the C-axis diminishes completely in 1.8cm thick Ho crystal. However,in high energy region, the depolarization can be small since the depolarization parameter D is inversely proportional to the neutron energy. We measured the energy dependence of neutron depolarization in the region from 0.1 eV to a few eV using a 4mm thick Ho single crystal and a polarized ^{3}He spin analyzer . We parametrize the measured depolarization as

P/P_{0} = exp(−2Dt)  (1)

and obtained

D = (0.104 ± 0.03eV · cm^{−1})/E_{n}  (2)

where P is the polarization of transmitted neutrons through the ferromagnetic Ho (4.2K), P_{0} is that through the normal Ho (135K) and t is the thickness of the Ho crystal (4mm). The results indicate that the depolarization can be ignored in our five-vector correlation experiment. As the configuration of our depolarization experiment is different from that at LosAlamos ,the direct comparison is not possible. However using a simple model for the depolarization mechanism, the two results look like disagree. To investigate this difference,we have measured neutron depolarization with the same crystal as the one used in LosAlamos. The preliminal result is consistent with the result of the LosAlamos group. It was also found that the neutron polarization recovers its value above a few tens eV. The analysis is now procceding . The difference between our crystal and the LosAlamos crystal is not understood. We also plan to measure the nuclear alignment by observing γ ray from an aligned Ho using our crystal.

(Ken-ichi OKUMURA)

References

[1] J.E.Koster,E.D.Davis,C.R.Gould,D.G.Haase,N.R.Roberson, L.W.Seagondoller,S.Wilburn,X.Zhu,Phys.Lett.B267,23(1991)

[2] D.G.Haase,J.D.Bowman,P.P.J.Delheij,C.M.Frankle,C.R.Gould, J.N.Knudson,J.E.Koster,G.E.Mitchell,S.Penttila,H.Postma, N.R.Roberson,S.J.Seestrom,S.H.Yoo,Phys.Rev.B46,11290(1992)